# LIVS: A Pluralistic Alignment Dataset for Inclusive Public Spaces

### v1

@@ Abstract
We introduce the Local Intersectional Visual Spaces (LIVS) dataset, a benchmark for multi-criteria alignment of text-to-image (T2I) models in inclusive urban planning. Developed through a two-year participatory process with 30 community organizations, LIVS encodes diverse spatial preferences across 634 initial concepts, consolidated into six core criteria—Accessibility, Safety, Comfort, Invitingness, Inclusivity, and Diversity—through 37,710 pairwise comparisons. Using Direct Preference Optimization (DPO) to fine-tune Stable Diffusion XL, we observed a measurable increase in alignment with community preferences, though a significant proportion of neutral ratings highlights the complexity of modeling intersectional needs. Additionally, as annotation volume increases, accuracy shifts further toward the DPO-tuned model, suggesting that larger-scale preference data enhances fine-tuning effectiveness. LIVS underscores the necessity of integrating context-specific, stakeholder-driven criteria into generative modeling and provides a resource for evaluating AI alignment methodologies across diverse socio-spatial contexts.
@@ 1 Introduction
Recent advances in text-to-image (T2I) generative modeling have significantly improved image quality and diversity (Bai et al., 2022; Podell et al., 2023; Zhang et al., 2024a). These developments can benefit communities by democratizing design processes in architecture, urban planning, and environmental visualization (Corner, 1999; Dubey et al., 2024). However, aligning T2I models with the specific needs of local communities remains an open challenge (Qadri et al., 2023; Kannen et al., 2024; Kirk et al., 2024), particularly when addressing subjective concepts such as inclusivity or safety.
@@
We collected 35,510 multi-criteria preference annotations, each covering one to three criteria, to fine-tune a Stable Diffusion XL model using Direct Preference Optimization (DPO) (Wallace et al., 2023; Rafailov et al., 2024). We then tested the fine-tuned model with 2,200 additional annotations, comparing it to the baseline model. In these comparisons, 700 favored the DPO model, 300 favored the baseline, and about 1,100 were neutral, highlighting the subjective and plural nature of alignment in this setting. Our experiments reveal how multi-criteria feedback can guide T2I models toward locally meaningful outputs, while illustrating that no single objective can fully capture the complexity of community priorities. Moreover, we observe that as the number of annotations increases, accuracy improves toward DPO, with criteria receiving more annotations showing a stronger preference for the fine-tuned model.
@@
  • We provide empirical evidence that DPO fine-tuning can modulate image generation according to multi-criteria feedback, revealing varying preferences across intersectional identities and a significant proportion of neutral responses, highlighting areas where preferences are balanced or where further refinement is needed to accommodate complex intersectional needs (Fan et al., 2023; Casper et al., 2023; Li et al., 2024; Rafailov et al., 2024).
@@ 2.1 Alignment of Generative Models
Studies on T2I alignment often focus on aesthetic preferences (Pressman et al., 2022; Kirstain et al., 2023; Wu et al., 2023) or content policy compliance. However, they typically assume a single, global notion of “goodness” or “suitability.” Pluralistic alignment, in contrast, recognizes that social values are heterogeneous and context-dependent (Turchin, 2019a; Kirk et al., 2024; Sorensen et al., 2024). Our work extends beyond purely global alignment by collecting specialized, multi-criteria annotations grounded in local, intersectional community knowledge for urban public space design.
Fewer efforts target image-based generative models compared to alignment in large language models (Bai et al., 2022; Huang et al., 2024b). Prior research has primarily utilized reward modeling and reinforcement learning from human feedback (Stiennon et al., 2022; Ouyang et al., 2022), with a focus on single-objective tasks such as helpfulness or factual correctness (Kirk et al., 2024; Jin et al., 2024). In contrast, our approach applies multi-criteria preference learning (Fan et al., 2023; Chakraborty et al., 2024) to T2I outputs within an urban planning context, capturing nuanced trade-offs among accessibility, safety, comfort, invitingness, inclusivity, and diversity.
@@ 2.2 Intersectionality and Local Knowledge
Intersectionality recognizes that individuals may experience multiple, overlapping forms of marginalization, affecting how they engage with public spaces and technology (Crenshaw, 1989; Costanza-Chock, 2020). In generative modeling, this perspective is often overlooked, with systems calibrated to an “average” user profile that can obscure the distinct needs of marginalized groups (Benjamin, 2019; Birhane, 2021; Gebru et al., 2021; Kirk et al., 2024; Murgia, 2024). For urban planning tasks, ignoring intersectionality risks overlooking critical insights related to accessibility, safety, or cultural expression. Integrating local knowledge adds further granularity, accounting for the historical, spatial, and communal context that global datasets typically lack (Fischer, 2000; Nekoto et al., 2020; Mohamed et al., 2020; D’Ignazio & Klein, 2020). While some studies acknowledge the importance of localized, context-specific input (Aroyo & Welty, 2015; Sloane, 2024; Sieber et al., 2024b), few systematically address intersectionality in T2I alignment. By weaving intersectional considerations into local community-driven annotations, our approach endeavors to reflect a broader range of perspectives and needs, moving beyond singular, one-size-fits-all criteria in generative modeling.
@@ 2.3 Visual Generative Modeling for Urban Spaces
Urban planning and design have long leveraged visualizations to communicate design objectives and gather feedback from stakeholders (Corner, 1999; Dubey et al., 2024; Guridi et al., 2024). T2I models offer the promise of more rapid prototyping and inclusive deliberation, especially when non-experts can directly prompt a model to generate conceptual designs of a plaza, park, or street (Rico Carranza et al., 2023; Guridi et al., 2024; Dubey et al., 2024; Agnew et al., 2024). Yet, generative models frequently default to learned global priors, potentially reproducing biases or neglecting local cultural markers (Hanna et al., 2024; Jin et al., 2024; Kirk et al., 2024; Sorensen et al., 2024). Our LIVS dataset is explicitly curated to capture local, intersectional preferences in a domain where the geometry, aesthetics, and sociocultural elements of a public space are all crucial (Jacobs, 1961; Gehl & Svarre, 2013; Mitrašinović & Mehta, 2021; Conitzer et al., 2024; Guridi et al., 2024; Dubey et al., 2024; Agnew et al., 2024). This approach aids in systematically evaluating how T2I alignment can be guided by multiple, sometimes conflicting, user-defined criteria.
@@ 3 Methodology: Building the LIVS Dataset
We employed a community-based participatory approach to integrate the local perspectives and experiences that are often absent in top-down, universal datasets (Sieber et al., 2024a; Kirk et al., 2024; Hosking et al., 2024). This methodology positions community members as collaborators, allowing for the identification of context-specific needs and priorities, such as nuanced understandings of accessibility and safety (Arnstein, 1969; Anttiroiko & de Jong, 2020). By involving diverse local organizations throughout the process, we aimed to capture intersectional viewpoints and mitigate the risk of imposing external definitions of inclusive design (IRCGM, 2018; Costanza-Chock, 2020). This approach also fosters mutual learning, wherein participants gain insights into AI technologies while researchers obtain domain-specific knowledge critical for aligning generative models with real-world public space requirements (Engeström, 2014).
@@ 3.1 Community Outreach and Engagement
We contacted 100 community organizations in a mid-sized city (Montreal, population $\sim$2 million) (Gouvernement du Canada, 2022). These organizations included neighborhood councils, disability-focused nonprofits, faith-based groups, youth advocacy networks, and other civic stakeholders. Our objective was to obtain a demographically diverse set of participants who frequently interact with local public spaces (IRCGM, 2018; Creswell & Creswell, 2022).
Figure 1: Distribution of Participants’ Self-Declared Demographics. This figure summarizes the demographic profiles (e.g., age, gender, race/ethnicity, and disability) of the individuals who participated in workshops and annotation activities.
@@
Over two years, we conducted multiple forms of engagement: eleven workshops, five batches of annotations, and 34 interviews. The process began in 2023 with three multi-stakeholder workshops aimed at defining what “equitable design” means for local public spaces. Participants from diverse backgrounds reviewed images of Montreal’s public spaces and discussed attributes they considered most important for inclusion.
  • Introductory Sessions (2 workshops, 25–35 participants each): Provided an introduction to AI and T2I technology. Participants shared open-ended feedback on the study, AI, and their experiences in local public spaces.
@@
  • Validation (1 workshop, 18 participants + 34 interviews): The 634 concepts were consolidated into 35 intermediate criteria through merging and semantic grouping. Participants then ranked and refined these, producing six final criteria based on perceived importance and local relevance.
@@
  • Prompting (1 workshop, 24 participants): Five groups, each composed of 2-3 citizens, a computer scientist, and an urban architect, collaboratively generated 440 prompts reflecting a variety of public-space scenarios and features. For instance, one group focused on designing prompts for pedestrian promenades with green spaces and safe-street initiatives in historical neighborhoods (see Appendix E for additional details).
  • Tutorial and Feedback (1 workshop, 20 participants): Using prompts from the previous workshop, we generated initial images with Stable Diffusion XL (Podell et al., 2023). Four groups, each comprising 3–4 citizens and an urban architect, tested the annotation platform and provided preliminary feedback on visual fidelity and representation before the larger-scale annotation phase (see Appendix C.3, Figure 11 for annotation platform).
@@
  • Annotations (5 batches, 18 participants): The annotation tasks were divided into five batches, each lasting two weeks and containing approximately 750 pairwise comparisons per participant, totaling 42,235 raw comparisons. In each task, two images were displayed side by side with three randomly selected criteria from the six. Annotators used a slider ranging from $-1$ (strong preference for the left image) to $+1$ (strong preference for the right image), with 0 indicating neutrality. An open-source annotation platform was developed to facilitate this process, featuring a user-friendly slider interface to accommodate diverse backgrounds (see Appendix D for additional details). A multi-stage data-cleaning process refined the dataset, resulting in 35,510 high-quality annotations (Prabhakaran et al., 2021).
@@ 3.2 Criteria Consolidation
The original 634 concepts spanned physical, social, and psychological attributes of public spaces (e.g., lighting, presence of diverse user groups, multilingual signage, seating). Through merging, voting, and iterative discussion, participants identified six core criteria (Figure 2 illustrates this process):
Accessibility: Physical and cognitive usability for people of all abilities, including ramps, tactile indicators, and clear signage.
Safety: Freedom from crime, hazards, or harassment, often reflected in well-lit areas, clear visibility, or protective barriers.
Comfort: Availability of amenities (e.g., seating, shade) and mitigation of environmental conditions (e.g., noise, temperature).
Invitingness: Features that encourage people to enter and remain, such as greenery, open layouts, or visible communal areas.
@@ 3.3 Prompt and Image Preparation
Figure 3: Word cloud depicting the frequency of various concepts within the 440 collected prompts. The size of each word reflects its prevalence, highlighting key themes such as public-space typologies, amenities, and contextual scenarios. This distribution underscores the diversity and comprehensiveness of the prompt dataset.
@@
For each image pair, participants were randomly assigned three of the six criteria and required to annotate at least one criterion. This modification, implemented following feedback from the evaluation workshop, ensured that each criterion had an equal opportunity to be evaluated. However, participants reported that certain criteria, such as Inclusivity, were more subjective, leading to fewer distinct preferences and a higher incidence of neutral or equal annotations. Consequently, while each criterion had an equal chance of being selected, some received fewer definitive annotations compared to others. This is reflected in Figure 4, where Inclusivity received the fewest annotations, whereas Comfort and Invitingness were annotated more frequently. Overall, the annotation process yielded 37,710 image-level comparisons across multiple criteria, amounting to approximately 113,130 annotations when divided per criterion. While the dataset size is modest by machine learning standards, the methodology prioritizes data quality, which helps mitigate limitations related to quantity.
@@ 4.1 Case Study I: Does Multi-Criteria DPO Improve Alignment?
We fine-tuned Stable Diffusion XL (SDXL; Podell et al. 2023) using DPO; Rafailov et al. 2024. We treated each pairwise comparison in LIVS as a binary “preferred” vs. “not preferred” signal, collapsing multi-criteria feedback via majority voting if annotators gave split ratings (e.g., preferring the left image for Safety but the right image for Inclusivity). Model outputs were then generated for a held-out set of prompts and compared against the SDXL baseline.

### v2

@@ Abstract
We introduce the Local Intersectional Visual Spaces (LIVS) dataset, a benchmark for multi-criteria alignment, developed through a two-year participatory process with 30 community organizations to support the pluralistic alignment of text-to-image (T2I) models in inclusive urban planning. The dataset encodes 37,710 pairwise comparisons across 13,462 images, structured along six criteria—Accessibility, Safety, Comfort, Invitingness, Inclusivity, and Diversity—derived from 634 community-defined concepts. Using Direct Preference Optimization (DPO), we fine-tune Stable Diffusion XL to reflect multi-criteria spatial preferences and evaluate the LIVS dataset and the fine-tuned model through four case studies: (1) DPO increases alignment with annotated preferences, particularly when annotation volume is high; (2) preference patterns vary across participant identities, underscoring the need for intersectional data; (3) human-authored prompts generate more distinctive visual outputs than LLM-generated ones, influencing annotation decisiveness; and (4) intersectional groups assign systematically different ratings across criteria, revealing the limitations of single-objective alignment. While DPO improves alignment under specific conditions, the prevalence of neutral ratings indicates that community values are heterogeneous and often ambiguous. LIVS provides a benchmark for developing T2I models that incorporate local, stakeholder-driven preferences, offering a foundation for context-aware alignment in spatial design.
@@ 1 Introduction
Recent advances in text-to-image (T2I) generative modeling have significantly improved image quality and diversity (Bai et al., 2022; Podell et al., 2023; Zhang et al., 2024a). These developments can benefit communities by making design processes more accessible—enabling broader engagement among non-expert stakeholders in architecture, urban planning, and environmental visualization. (Corner, 1999; Pallasmaa, 2011; Dubey et al., 2024). However, aligning T2I models with the specific needs of local communities remains an open challenge (Qadri et al., 2023; Kannen et al., 2024; Kirk et al., 2024), particularly when addressing subjective concepts such as inclusivity or safety.
@@
We collected 35,510 multi-criteria preference annotations, each covering one to three criteria, to fine-tune a Stable Diffusion XL model using Direct Preference Optimization (DPO) (Wallace et al., 2023; Rafailov et al., 2024). We then tested the fine-tuned model with 2,200 additional annotations, comparing it to the baseline model. In these comparisons, 700 favored the DPO model, 300 favored the baseline, and about 1,100 were neutral, indicating that existing alignment methods are insufficient to capture diverse, subjective preferences and highlighting the need for more nuanced alignment approaches. Our experiments demonstrate that multi-criteria feedback can steer T2I models toward locally meaningful outputs, while also showing that no single alignment objective can account for the complexity of community priorities. Additional case studies demonstrate that preference patterns differ across participant identities, that human-authored prompts result in more visually distinct outputs than those generated by language models, and that intersectional groups provide systematically different ratings across criteria, further illustrating the limitations of single-objective alignment frameworks.
@@
  • We provide empirical evidence that DPO fine-tuning can modulate image generation according to multi-criteria feedback, yet the prevalence of neutral responses indicates the continued insufficiency of current alignment methods to capture complex intersectional preferences, underscoring the need for approaches capable of accommodating nuanced, diverse user needs. (Fan et al., 2023; Casper et al., 2023; Li et al., 2024; Rafailov et al., 2024).
@@ 2.1 Alignment of Generative Models
Studies on T2I alignment often focus on aesthetic preferences (Pressman et al., 2022; Kirstain et al., 2023; Wu et al., 2023) or content policy compliance. However, they typically assume a single, global notion of “goodness” or “suitability.” Pluralistic alignment, in contrast, recognizes that social values are heterogeneous and context-dependent (Turchin, 2019a; Sorensen et al., 2024). Our work extends beyond purely global alignment by collecting specialized, multi-criteria annotations grounded in local, intersectional community knowledge for urban public space design.
Fewer efforts target image-based generative models compared to alignment in large language models (Bai et al., 2022; Huang et al., 2024b). Prior research has primarily utilized reward modeling and reinforcement learning from human feedback (Stiennon et al., 2022; Ouyang et al., 2022), with a focus on single-objective tasks such as helpfulness or factual correctness (Kirk et al., 2024; Jin et al., 2024). In contrast, our approach leverages multi-criteria annotations to inform alignment in T2I outputs within an urban planning context, though these signals are ultimately aggregated into a single binary label for model training via majority voting.
@@ 2.2 Intersectionality and Local Knowledge
Intersectionality recognizes that individuals may experience multiple, overlapping forms of marginalization, affecting how they engage with public spaces and technology (Crenshaw, 1989; Costanza-Chock, 2020). In generative modeling, this perspective is often overlooked, with systems calibrated to an “average” user profile that can obscure the distinct needs of marginalized groups (Benjamin, 2019; Gebru et al., 2021; Kirk et al., 2024; Murgia, 2024). For urban planning tasks, ignoring intersectionality risks overlooking critical insights related to accessibility, safety, or cultural expression. Integrating local knowledge adds further granularity, accounting for the historical, spatial, and communal context that global datasets typically lack (Fischer, 2000; Nekoto et al., 2020; Mohamed et al., 2020; D’Ignazio & Klein, 2020). While some studies acknowledge the importance of localized, context-specific input (Aroyo & Welty, 2015; Sloane, 2024; Sieber et al., 2024b), few systematically address intersectionality in T2I alignment. By weaving intersectional considerations into local community-driven annotations, our approach endeavors to reflect a broader range of perspectives and needs, moving beyond singular, one-size-fits-all criteria in generative modeling.
@@ 2.3 Visual Generative Modeling for Urban Spaces
Urban planning and design have long leveraged visualizations to communicate design objectives and gather feedback from stakeholders (Corner, 1999; Dubey et al., 2024; Guridi et al., 2024). T2I models offer the promise of more rapid prototyping and inclusive deliberation, especially when non-experts can directly prompt a model to generate conceptual designs of a plaza, park, or street (Guridi et al., 2024; Dubey et al., 2024). Yet, generative models frequently default to learned global priors, potentially reproducing biases or neglecting local cultural markers (Hanna et al., 2024; Jin et al., 2024; Kirk et al., 2024; Sorensen et al., 2024). Our LIVS dataset is explicitly curated to capture local, intersectional preferences in a domain where the geometry, aesthetics, and sociocultural elements of a public space are all crucial (Jacobs, 1961; Gehl & Svarre, 2013; Mitrašinović & Mehta, 2021; Conitzer et al., 2024; Guridi et al., 2024; Dubey et al., 2024). This approach aids in systematically evaluating how T2I alignment can be guided by multiple, sometimes conflicting, user-defined criteria.
@@ 3 Methodology: Building the LIVS Dataset
We employed a community-based participatory approach to integrate the local perspectives and experiences that are often absent in top-down, universal datasets (Sieber et al., 2024a; Hosking et al., 2024). This methodology positions community members as collaborators, allowing for the identification of context-specific needs and priorities, such as nuanced understandings of accessibility and safety (Arnstein, 1969; Anttiroiko & de Jong, 2020; Mushkani et al., 2025a). By involving diverse local organizations throughout the process, we aimed to capture intersectional viewpoints and mitigate the risk of imposing external definitions of inclusive design (Costanza-Chock, 2020). This approach also fosters mutual learning, wherein participants gain insights into AI technologies while researchers obtain domain-specific knowledge critical for aligning generative models with real-world public space requirements (Engeström, 2014).
@@ 3.1 Community Outreach and Engagement
We contacted 100 community organizations in a mid-sized city (Montreal, population $\sim$2 million) (Gouvernement du Canada, 2022). These organizations included neighborhood councils, disability-focused nonprofits, faith-based groups, youth advocacy networks, and other local civic stakeholders. Our objective was to obtain a demographically and experientially diverse set of participants who frequently interact with local public spaces (IRCGM, 2018; Creswell & Creswell, 2022).
Participatory Action Research (PAR).
In line with the principles of PAR, our community-based approach centers on iterative, collaborative inquiry and reciprocal learning throughout the dataset development process (Israel et al., 1998; Cornish et al., 2023). By involving local organizations as active co-researchers, we ensured that the framing of inclusion, safety, and other design criteria emerged from lived experiences rather than external prescriptions. This iterative feedback loop aligns with PAR’s emphasis on collective problem-solving and empowerment, as participants guided each stage of data collection and model evaluation while gaining familiarity with T2I technology and its potential applications in urban contexts.
@@
Over two years, we conducted multiple forms of engagement: eleven workshops, five batches of annotations, and 34 interviews. The process began in 2023 with three multi-stakeholder workshops aimed at collaboratively defining what “equitable design” means for local public spaces. Participants from diverse backgrounds reviewed images of Montreal’s public spaces and discussed attributes they considered most important for inclusion. Figure 1 shows the distribution of participants’ self-declared demographics.
  • Introductory Sessions (2 workshops, 25–35 participants each): Provided an introduction to AI and T2I technology. Participants also shared open-ended feedback on the study, AI, and their experiences in local public spaces.
@@
  • Validation (1 workshop, 18 participants + 34 interviews): The 634 concepts were consolidated into 35 intermediate criteria through merging and semantic grouping. Participants then ranked and refined these, producing six final criteria—Accessibility, Safety, Comfort, Invitingness, Inclusivity, and Diversity—based on importance and local relevance.
Figure 1: Distribution of participants’ self-declared demographics. This figure summarizes the demographic profiles (e.g., age, gender, race/ethnicity, and disability) of the individuals who participated in workshops and annotation activities (Mushkani et al., 2025b).
@@
  • Prompting (1 workshop, 24 participants): Five groups, each composed of 2-3 citizens, a computer scientist, and an urban architect, collaboratively generated 440 prompts reflecting a variety of public-space scenarios and features. For instance, one group focused on designing prompts for pedestrian promenades with green spaces and safe-street initiatives in historical neighborhoods (see Appendix E for additional details). While refined prompting techniques can shape generative outputs, universal keywords alone may fail to capture local sociocultural and historical contexts (Anttiroiko & de Jong, 2020; Beebeejaun, 2016; Talen, 2012). The LIVS dataset was developed to incorporate granular, community-generated perspectives on accessibility, safety, and inclusivity. By embedding localized knowledge, we aim to reduce the risk of producing uniform design standards that overlook intersectional needs (Low, 2020; Madanipour, 2010; McAndrews et al., 2023).
  • Tutorial and Feedback (1 workshop, 20 participants): Using prompts from the previous workshop, we generated initial images with Stable Diffusion XL (Podell et al., 2023). Four groups, each comprising 3–4 citizens and an urban architect, tested the annotation platform and provided feedback on visual fidelity and representation before the larger-scale annotation phase (see Appendix C.3, Figure 12 for annotation platform). To optimize data quality and minimize cognitive fatigue, we randomly presented three of the six criteria in each comparison. During pilot trials, participants found evaluating all six criteria difficult, which reduced their sense of progress and visual engagement. The multiple rating elements also constrained image size, making it harder to assess spatial details. Focusing on three criteria enabled more meaningful engagement. Over successive annotation batches, we ensured that all six criteria were robustly evaluated.
@@
  • Annotations (5 batches, 18 participants): The annotation tasks were divided into five batches, each lasting two weeks and containing approximately 750 pairwise comparisons per participant, totaling 42,235 raw comparisons. In each task, two images were displayed side by side with three randomly selected criteria from the six. Annotators used a slider ranging from $-1$ (strong preference for the left image) to $+1$ (strong preference for the right image), with 0 indicating neutrality. An open-source annotation platform was developed to facilitate this process, featuring a user-friendly slider interface to accommodate diverse backgrounds (see Appendix D for additional details). A multi-stage data-cleaning process refined the dataset, including removal of quality-control annotations and incomplete submissions. The resulting dataset comprises 35,510 high-quality annotations.
@@ 3.2 Criteria Consolidation
The original 634 concepts spanned physical, social, and psychological attributes of public spaces (e.g., lighting, presence of diverse user groups, multilingual signage, seating). Through collaborative merging, collective voting, and iterative discussion, participants identified six core criteria (Figure 2 illustrates this process):
Accessibility: Physical and cognitive usability for people of all abilities, including elevators, ramps, tactile indicators, and clear signage.
Safety: Freedom from crime, hazards, or harassment, often reflected in well-lit areas, clear visibility, active presence, or protective barriers.
Comfort: Availability of amenities (e.g., seating, shade) and mitigation of environmental conditions (e.g., noise, temperature), including cleanliness and spatial shelter.
Invitingness: Features that encourage people to enter and remain, such as greenery, open layouts, welcoming signage, or visible communal areas.
@@ 3.3 Prompt and Image Preparation
Figure 3: Word cloud depicting the frequency of various concepts within the 440 collected prompts. The size of each word reflects its prevalence, highlighting key themes such as public-space typologies, amenities, and contextual use scenarios. This distribution underscores the diversity and contextual comprehensiveness of the prompt dataset.
@@
For each image pair, participants were randomly assigned three of the six criteria and required to annotate at least one criterion. This modification, implemented following feedback from the evaluation workshop, ensured that each criterion had an equal opportunity to be evaluated. However, participants reported that certain criteria, such as Inclusivity, were more subjective, leading to fewer distinct preferences and a higher incidence of neutral or equal annotations. Consequently, while each criterion had an equal chance of being selected, some received fewer definitive annotations compared to others. This is reflected in Figure 4, where Inclusivity received the fewest annotations, whereas Comfort and Invitingness were annotated more frequently. Overall, the annotation process yielded 37,710 image-level comparisons, amounting to approximately 113,130 criterion-level annotations. Though modest in size by machine learning standards, the dataset prioritizes quality over volume.
@@ 4.1 Case Study I: Does Multi-Criteria DPO Improve Alignment?
We fine-tuned Stable Diffusion XL (SDXL; Podell et al. 2023) using Direct Preference Optimization (DPO; Rafailov et al. 2024). Each pairwise comparison in LIVS was treated as a binary “preferred” vs. “not preferred” signal. When annotators provided split ratings across criteria (e.g., preferring the left image for Safety but the right image for Inclusivity), we collapsed the feedback via majority voting. This procedure trained a single model on all criteria and evaluated its performance on each criterion individually. Model outputs were then generated for a held-out set of prompts and compared against the SDXL baseline.